Radius of capillary is 2mm. A liquid of weight 6.28 x 10^-4N may remain in the capillary, then the surface tension of liquid will be?


1
Expert's answer
2021-09-09T10:49:00-0400

"γ = \frac{F}{l}"

"\text{The surface tension forces F are numerically equal to the weight }"

"\text{of the liquid in the capillary.Since the liquid in the capillary}"

"\text{is in equilibrium.}"

"F = 6.28*10^{-4}N"

"l =2\pi r"

"r = 2mm = 2*10^{-3}m"

"γ = \frac{F}{l}= \frac{6.28*10^{-4}}{2*\pi*2*10^{-3}}\approx0.05\frac{N}{m}"


"\text{Answer: }0.05\frac{N}{m}"